OpenStudy (anonymous):

If vector a=-8(vector b) and vector c=7(vector b) and what is angle between a and c ??

OpenStudy (amistre64):

cos(t) = a*c/|a||c|

OpenStudy (amistre64):

b*b = |b|^2

OpenStudy (amistre64):

<-8xb, -8yb> < 7xb, 7yb> -------------- -xb + -yb = -1<b>

OpenStudy (amistre64):

thats messed up lol

OpenStudy (anonymous):

yeah.. i didnt really understand that..

OpenStudy (anonymous):

what do u mean in ur 3rd reply..??

OpenStudy (amistre64):

a*c = -8<xb,yb> * 7<xb,yb>

OpenStudy (amistre64):

-56x^2b + -56 y^b is the only values we can get from it right?

OpenStudy (amistre64):

-56|b|^2 is the top value

OpenStudy (amistre64):

the bottom is -8|b| * 7|b| = -56|b|^2 cos(t) = 1

OpenStudy (amistre64):

t = 90 begrees right?

OpenStudy (amistre64):

so hard to type in the dark lol

OpenStudy (amistre64):

or it might just be 180 degrees; since a and c are scalars of b and they are facing inopposite directions.

OpenStudy (anonymous):

I worked it similarly. I don't know if we got same place. \[-8b.-7b =\left| a \right|\left| c \right|\cos \theta\] \[-8b.-7b =8b7b \cos \theta\] \[-1=\cos \theta\] \[\cos^{-1} -1=\theta\] \[\theta = 3.14 (radians)\]

OpenStudy (anonymous):

I didn't dot the left hand side. May be that is a mistake that needs to be cleaned up. I should go to sleep now.